Please show how Write the point-slope form of the equation of the line passing through the points
(1,4) and (3,6).



Answer by robertb(5830)

The slope is (6-4)/(3-1) = 2/2 = 1.
Then the point slope form is either y - 4 = 1*(x - 1), or y - 6 = 1*(x - 3).